[clas_members] [New Logentry] RC Daily Summary

daria at jlab.org daria at jlab.org
Fri Sep 9 13:15:01 EDT 2022

Logentry Text:

Hello everyone!

We have been running more or less steadily, with minor glitches since yesterday. 

Hall C did a BCM calibration in the early afternoon yesterday, taking our beam for 1h20. They have stopped running at 8am this morning and we are now the priority hall for the coming week. 

We took some very nice data yesterday during the day. 

In the evening, Maxwell found that the remote control for the NMR frequency wasn't working, so the frequency couldn't be fine-tuned to give maximum polarisation. We ran with slightly reduced polarisation, with it dropping into the twenties by this morning. Maxwell and James did a rapid access to debug (about 1h30), the controls are working now but need to be driven manually from the counting house: 
This is an interim fix, until a more in-depth debugging can take place --- oh, call from Maxwell, he says he thinks he can fix it in 5min, will be making a rapid access for this shortly!

Yesterday evening there was a DAQ crash (run 16907) and a chunk of FTOF S6 TDCs vanishing (run 16908, should probably be junked): https://logbooks.jlab.org/entry/4039345

We took this as a sign to change half-wave plate. 

In the middle of the night an MCC IOC controller went, without any apparent alarms for them -- some beamline devices started showing as disconnected for us: https://logbooks.jlab.org/entry/4039454
Loss of beam (to us and Hall C): about 30min for them to find and reboot it.

Now we're back up, taking data and crossing fingers and toes for some more smooth sailing. 

Your generally quite happy RC :-)

**** Stats since the start of ND3 with ELMO ****
Half-wave plate OUT, target polarisation POSITIVE:

16902: 100.4 M (pinned)
16903:  102 M 
16904: 100.1 M (pinned)
16905: 121.8 M
16907: 42.6 M (DAQ crash at end)
16908: 37.4 M (presumed junk! missing part of FTOF)

TOTAL triggers from "good" runs: 467 M

Half-wave plate IN, target polarisation POSITIVE:

16910: 91.1 M
16911: 101.6 M
16912: 71.8 M (beamline devices disconnected at the end)
16913: 86.4 M
16918: currently at 39.5M, still running... (pinned)

TOTAL triggers from "good" runs so far: 390 M 
(but target polarisation was lower than for the half-wave plate OUT data, as there was a problem with NMR controls...)


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4039575
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/clas_members/attachments/20220909/494dd6ae/attachment.html>

More information about the clas_members mailing list